# Dynamic borrowing from historical controls via the synthetic prior with covariates in randomized clinical trials

### v1

@@ 1 Introduction
From a regulatory perspective, historical controls have been permitted in confirmatory trials primarily in rare and pediatric diseases as well as in devices (Ghadessi et al. 2020). However, the regulatory threshold for their use is lower in non-confirmatory settings where there is less demand for conservative Type I error guarantees (U.S. Food and Drug Administration 2019). In order to use historical controls effectively, statistical models are critical due to the challenge in reconciling historical data and concurrent control data. Ideally, historical data that are more “similar” should be borrowed from more to aid statistical inference. The main questions are how to measure the similarity and how to “borrow” based on this measure.
@@
Popular statistical methods for leveraging historical controls include propensity score approaches, which typically match or weight historical and current patients based on covariates (Lim et al. 2018, Lin et al. 2018, Chen et al. 2022), as well as Bayesian modeling strategies including meta-analytic priors such as MAP and RMAP (Neuenschwander et al. 2010; Schmidli et al. 2014), power priors (Chen & Ibrahim 2000), commensurate priors (Hobbs et al. 2012), multisource exchangeability models (Kaizer et al. 2018), and LEAP (Alt et al., 2024), which tend to borrow based primarily on similarities in response rates between the historical and new data. A barrier to using propensity score methods is that they require rich patient-level data, which is often unavailable (e.g. when such data are owned by a competing developer), and that researchers must both have and select a sufficient set of covariates to control bias. While the Bayesian methods are attractive strategies to dynamically adapt the degree of historical borrowing, they typically assume a single mechanism of borrowing: either that the historical and new data have similar unadjusted response rates, or that they have similar response-covariate regression functions. As we show, this can lead to a loss of robustness or efficiency in scenarios violating the assumed borrowing mechanism.
@@
We propose a new model called SPx, standing for “synthetic prior with covariates,” to sharpen inferences about a new trial’s control group response rate by borrowing from historical data. The main idea of SPx is to use a Bayesian expert system (Spiegelhalter et al. 1993) to merge different mechanisms of information sharing between the historical and new trial data. This strategy allows the model to efficiently tailor its inferences to different settings where distinct mechanisms of historical borrowing (or none) are appropriate. The model can be used simply for the analysis of a completed trial, but we also discuss how it can be embedded in an adaptive trial design to reduce the needed control arm sample size.
@@
Similar to popular methods like MAP and RMAP (Neuenschwander et al. 2010; Schmidli et al. 2014), SPx requires only summary statistics from the historical trials, not patient-level data. Such trial-level information is routinely reported in publications and press releases, and includes sample size, response rate, eligibility criteria, and average patient demographics and pre-trial clinical measures. This means that researchers using SPx may potentially draw from many more historical trials than if they were to use methods that require patient-level data. While patient-level are ideal to use for historical borrowing, they are often unavailable due to ethical and confidentiality reasons. Nevertheless, when patient data are available, the SPx model can be modified easily to accommodate the new data structure. We discuss this point briefly at the end of the paper.
@@ 2.1 Basic Data Setting and Related Models
Data structure. Following Schmidli et al. (2014), we consider summary statistics of patient-level data in historical clinical trials. Specifically, denote the data by $(y_{h},n_{h},\mathbf{x}_{h})$ for trials $h=1,\dots,H,H+1$. Trials $1,\dots,H$ are the historical trials and trial $H+1$ is the new trial, $y_{h}$ is the number of responders in trial $h$’s control group (from a binary endpoint), $n_{h}$ is the number of control patients in trial $h$, and $\mathbf{x}_{h}$ is a $(p+1)$-dimensional vector containing $p$ group-level covariates for trial $h$’s control group as well as an intercept.
@@
Meta-analytic priors. The MAP, or meta-analytic predictive prior (Neuenschwander et al. 2010) is exchangeable, and models the logit response rates $\theta_{h}:=\text{logit}(\psi_{h})$ of all the trials (historical and new) with a common normal prior distribution:
@@
The prior mean $\mu$ and variance $\tau^{2}$ recieve hyperpriors to complete the hierarchical model specification. Although Neuenschwander et al. (2010) emphasize the predictive interpretation of the MAP prior, here we focus on its hierarchical interpretation to illustrate how it induces borrowing through shrinkage. In MAP the variance parameter $\tau^{2}$ controls the degree of historical borrowing: if $\tau^{2}$ is small then the posterior will shrink $\theta_{H+1}$ strongly towards the historical data, and if $\tau^{2}$ is large then historical borrowing will be curtailed. Thus the hyperprior for $\tau^{2}$ is crucial to the performance of the MAP method, and in the original work Neuenschwander et al. (2010) recommended sensitivity analysis for this part of the prior. In contrast, MAP’s performance does not depend as heavily on the hyperprior for $\mu$ as long as it is not unreasonably concentrated. Often a noninformative uniform prior is used for $\mu$. The MAP model’s popularity stems from its simple, familiar random effects model and good performance when the historical control rates are largely similar to the new trial’s control rate. However, when the historical data are misleading the MAP approach often continues to borrow too heavily and thus lacks robustness.
Due to the MAP approach’s strong reliance on the historical data’s relevance to avoid bias and inflated Type I error, a Robust MAP (RMAP) model was also proposed (Schmidli et al. 2014). This extension is a combination of the MAP prior and an independent prior that involves no borrowing from historical data, given by
@@
Commensurate priors. Another approach to historical borrowing is the commensurate prior (Hobbs et al. 2012), which explicitly models the difference between the historical and new trials (Pocock 1976). In our setting it first assumes that all historical rates are equal and then specifies the joint prior on the historical and new rates through a marginal prior on the historical rate and a “commensurate” prior on the new rate given the historical one:
@@
Thus the (logit) difference between the new trial’s rate and the historical rate is modeled as $N(0,\sigma^{2})$, where the variance $\sigma^{2}$ is the key parameter controlling the amount of historical borrowing. Despite that assuming homogeneity of the response rates in the historical trials may be an oversimplification, this conditional specification provides a different mechanism to control borrowing than in the meta-analytic approaches. In past work (Hobbs et al. 2012) commensurate priors have been designed to work with patient-level data and covariates.
@@ 2.2 The SPx Prior
A key distinction with RMAP is that SPx uses submodels that are well-separated in the sense that they put most of their prior mass on separate regions of the parameter space for $\left(\theta_{1},\dots,\theta_{H+1}\right)$. This feature is motivated by the literature on nonlocal priors (Johnson & Rossel 2010). As we later discuss, it leads to posterior inferences about the mechanism and degree of historical borrowing that adapt more quickly to signals in the data. We now describe each component of the SPx prior in detail.
@@
$\theta_{H+1}$ $=z_{hist}\underbrace{\theta_{hist}}_{\text{Expert 1}}+z_{reg}\underbrace{\theta_{reg}}_{\text{Expert 2}}+z_{ind}\underbrace{\theta_{ind}}_{\text{Expert 3}}.$
Here the expert system, or model averaging, prior is obtained by assuming the Categorical prior $(z_{hist},z_{reg},z_{ind})\sim Cat(p_{hist},p_{reg},p_{ind}),$ meaning that the $z_{k}$ are binary random variables satisfying $z_{hist}+z_{reg}+z_{ind}=1$, and $\theta_{H+1}$ is drawn from model $k\in\{hist,reg,ind\}$ with prior probability $p_{k}$. Priors for each expert, i.e. $\theta_{hist},\theta_{ref},\text{and }\theta_{ind}$, are discussed next.
@@
$w_{h}\propto\left(0.5\right)^{\frac{|\tilde{\pi}_{h}-\tilde{\pi}_{H+1}|}{0.05}}$
with $\tilde{\pi}_{h}:=\text{logit}^{-1}(\bm{\beta}^{T}\mathbf{x}_{h})$. This choice encourages more borrowing from historical trials with covariates closer to $\mathbf{x}_{H+1}$, to the extent that the covariates predict response rates. We tested other functional forms for the weights and found they typically do not affect posterior inferences for $\hat{\theta}_{H+1}$ greatly (especially when $H$ is not very small).
The prior on $\theta_{reg}$, the regression expert, assumes that the new trial follows the same covariate-response relationship seen in the historical trials. This occurs because the same coefficient $\bm{\beta}$ is used in the regression for $\theta_{h}$ ($h=1,\dots,H$) in (1) and for $\theta_{reg}$. Another key point is that the scale depends on the same $\tau^{2}$ indicating how successful the regression is at predicting the historical response rates, but modified by a constant $c$ for reasons we discuss below along with the hyperpriors. At the point $\bm{\beta}=\bm{0}$ the regression expert is equivalent to the classic MAP model.
Lastly, $\theta_{ind}$ is the independent or no-borrowing expert. It is included for robustness, and uses a standard Jeffreys prior.
@@
$y_{h}|\mathbf{x}_{h},\psi_{h}$ $\overset{iid}{\sim}Bin(n_{h},\psi_{h}),\hskip 30.00005pth=1,\dots,H,H+1$
$\theta_{h}$ $:=\text{logit}(\psi_{h}),\hskip 30.00005pth=1,\dots,H,H+1$
@@
$\theta_{H+1}$ $=z_{hist}\theta_{hist}+z_{reg}\theta_{reg}+z_{ind}\theta_{ind},$
$\theta_{hist}|\mu_{hist},\sigma^{2}$ $\sim N(\mu_{hist},\sigma^{2}),\hskip 30.00005pt\mu_{hist}:=\sum_{h=1}^{H}w_{h}\theta_{h},$
$w_{h}$ $\propto\left(0.5\right)^{\frac{|\tilde{\pi}_{h}-\tilde{\pi}_{H+1}|}{0.05}},\hskip 30.00005pt\tilde{\pi}_{h}:=\text{logit}^{-1}(\bm{\beta}^{T}\mathbf{x}_{h})$
@@
$\text{logit}^{-1}(\theta_{ind})$ $\sim Beta(0.5,0.5),$
$(z_{hist},z_{reg},z_{ind})$ $\sim Cat(p_{hist},p_{reg},p_{ind}).$
@@
The prior model probabilities $(p_{hist},p_{reg},p_{ind})$ and the hyperpriors for the variances $\sigma^{2}$ and $\tau^{2}$ (and to a lesser degree $\bm{\beta}$) are important because they impact the posterior weighting of the different borrowing mechanisms in SPx. To see how, note that the posterior for $\theta_{H+1}$ is the average of the posteriors under each expert, weighting by the posterior probability that each is “correct”:
$p(\theta_{H+1}|D)=\sum_{k\in\{hist,reg,ind\}}p(\theta_{k}|z_{k}=1,D)\cdot p(z_{k}=1|D),$ (2)
where $D$ denotes the complete data $(y_{h},n_{h},\mathbf{x}_{h})_{h=1,\dots,H+1}$. The posterior weights in (2), $p^{*}_{k}:=p(z_{k}=1|D)$ for $k\in\{hist,reg,ind\}$, may be written as
$p^{*}_{k}\propto p_{k}\cdot p(D|z_{k}=1)$
where $p(D|z_{k}=1)=\int_{\mathbb{R}^{H+1}}p(D|\theta_{1},\dots,\theta_{H+1})\cdot p(\theta_{1},\dots,\theta_{H+1}|z_{k}=1)\hskip 3.0pt\mathrm{d}^{H+1}(\theta_{1},\dots,\theta_{H+1})$ is the marginal likelihood, or evidence, of expert $k$. It is important to recognize that each expert’s hyperprior affects their marginal prior on the logit rates, $p(\theta_{1},\dots,\theta_{H+1}|z_{k}=1)$. This in turn can greatly influence the expert’s marginal likelihood and posterior model probability, changing the behavior of SPx overall.
@@
where $TCauchy(m,s,I)$ represents a Cauchy distribution with location $m$ and scale $s$ that has been truncated to the interval $I$. While $\tau$ has a larger scale than $\sigma$, recall that the (conditional) prior variance of the regression submodel is $c\tau^{2}=\nicefrac{{1}}{{25}}\cdot\tau^{2}$, not $\tau^{2}$, so the regression submodel makes similarly strong prior predictions as the direct historical borrowing one. The prior on the regression coefficients, which does not unduly affect inference, is
@@
following from Gelman et al. (2008).
This choice of prior underlies SPx’s novel strategy to dynamically determine how much historical borrowing is appropriate. Combining (a) and (b) makes SPx more robust to irrelevant historical data as needed in equation (2) primarily by increasing $p^{*}_{ind}$, the posterior probability of the no-borrowing submodel, and only to a much lesser degree by making the posterior inferences $p(\theta_{k}|z_{k}=1,D)$ of the borrowing submodels more conservative. In further detail, (a) makes the borrowing submodels’ marginal likelihoods decrease more quickly as disagreement between the new and historical data grows, and (b) accounts for the fact that the no-borrowing submodel makes less confident predictions overall and thus may have a relatively lower marginal likelihood even when it should be favored. While similar robustness might be sought by giving the borrowing submodels more diffuse priors, this can come at the cost of weaker borrowing when the historical data are actually relevant. We discuss these points further in the supporting information.
@@
Combined with the SPx prior for the new trial’s control rate this induces a prior on the treatment effect, which we define as the difference $\delta:=\psi_{trt}-\psi_{H+1}$ (although other comparisons, such as relative risk, could just as easily be used).

### v2

@@ 1 Introduction
From a regulatory perspective, historical controls have been permitted in confirmatory trials primarily in rare and pediatric diseases as well as in devices [Ghadessi et al. 2020]. However, the regulatory threshold for their use is lower in non-confirmatory settings where there is less demand for conservative Type I error guarantees [U.S. Food and Drug Administration 2019]. In order to use historical controls effectively, statistical models are critical due to the challenge in reconciling historical data and concurrent control data. Ideally, historical data that are more “similar” should be borrowed from more to aid statistical inference. The main questions are how to measure the similarity and how to “borrow” based on this measure.
@@
Popular statistical methods for leveraging historical controls include propensity score approaches, which typically match or weight historical and current patients based on covariates [Lim et al. 2018, Lin et al. 2018, Chen et al. 2022], as well as Bayesian modeling strategies including meta-analytic priors such as MAP and RMAP [Neuenschwander et al. 2010, Schmidli et al. 2014], power priors [Chen and Ibrahim 2000], commensurate priors [Hobbs et al. 2012], multisource exchangeability models [Kaizer et al. 2018], and LEAP [Alt et al. 2024], which tend to borrow based primarily on similarities in response rates between the historical and new data. A barrier to using propensity score methods is that they require rich patient-level data, which is often unavailable (e.g. when such data are owned by a competing developer), and that researchers must both have and select a sufficient set of covariates to control bias. While the Bayesian methods are attractive strategies to dynamically adapt the degree of historical borrowing, they typically assume a single mechanism of borrowing: either that the historical and new data have similar unadjusted response rates, or that they have similar response-covariate regression functions. As we show, this can lead to a loss of robustness or efficiency in scenarios violating the assumed borrowing mechanism.
@@
We propose a new model called SPx, standing for “synthetic prior with covariates,” to sharpen inferences about a new trial’s control group response rate by borrowing from historical data. The main idea of SPx is to use a Bayesian expert system [Spiegelhalter et al. 1993] to merge different mechanisms of information sharing between the historical and new trial data. This strategy allows the model to efficiently tailor its inferences to different settings where distinct mechanisms of historical borrowing (or none) are appropriate. The model can be used simply for the analysis of a completed trial, but we also discuss how it can be embedded in an adaptive trial design to reduce the needed control arm sample size.
@@
Similar to popular methods like MAP and RMAP [Neuenschwander et al. 2010, Schmidli et al. 2014], SPx requires only summary statistics from the historical trials, not patient-level data. Such trial-level information is routinely reported in publications and press releases, and includes sample size, response rate, eligibility criteria, and average patient demographics and pre-trial clinical measures. This means that researchers using SPx may potentially draw from many more historical trials than if they were to use methods that require patient-level data. While patient-level are ideal to use for historical borrowing, they are often unavailable due to ethical and confidentiality reasons. Nevertheless, when patient data are available, the SPx model can be modified easily to accommodate the new data structure. We discuss this point briefly at the end of the paper.
@@ 2.1 Basic Data Setting and Related Models
[…]
@@
Meta-analytic priors. The MAP, or meta-analytic predictive prior [Neuenschwander et al. 2010] is exchangeable, and models the logit response rates $\theta_{h}:=\text{logit}(\psi_{h})$ of all the trials (historical and new) with a common normal prior distribution:
@@
The prior mean $\mu$ and variance $\tau^{2}$ recieve hyperpriors to complete the hierarchical model specification. Although Neuenschwander et al. [2010] emphasize the predictive interpretation of the MAP prior, here we focus on its hierarchical interpretation to illustrate how it induces borrowing through shrinkage. In MAP the variance parameter $\tau^{2}$ controls the degree of historical borrowing: if $\tau^{2}$ is small then the posterior will shrink $\theta_{H+1}$ strongly towards the historical data, and if $\tau^{2}$ is large then historical borrowing will be curtailed. Thus the hyperprior for $\tau^{2}$ is crucial to the performance of the MAP method, and in the original work Neuenschwander et al. [2010] recommended sensitivity analysis for this part of the prior. In contrast, MAP’s performance does not depend as heavily on the hyperprior for $\mu$ as long as it is not unreasonably concentrated. Often a noninformative uniform prior is used for $\mu$. The MAP model’s popularity stems from its simple, familiar random effects model and good performance when the historical control rates are largely similar to the new trial’s control rate. However, when the historical data are misleading the MAP approach often continues to borrow too heavily and thus lacks robustness.
Due to the MAP approach’s strong reliance on the historical data’s relevance to avoid bias and inflated Type I error, a Robust MAP (RMAP) model was also proposed [Schmidli et al. 2014]. This extension is a combination of the MAP prior and an independent prior that involves no borrowing from historical data, given by
@@
Commensurate priors. Another approach to historical borrowing is the commensurate prior [Hobbs et al. 2012; 2013], which explicitly models the difference between the historical and new trials [Pocock 1976]. In our setting it first assumes that all historical rates are equal and then specifies the joint prior on the historical and new rates through a marginal prior on the historical rate and a “commensurate” prior on the new rate given the historical one:
@@
Thus the (logit) difference between the new trial’s rate and the historical rate is modeled as $N(0,\sigma^{2})$, where the variance $\sigma^{2}$ is the key parameter controlling the amount of historical borrowing. Despite that assuming homogeneity of the response rates in the historical trials may be an oversimplification, this conditional specification provides a different mechanism to control borrowing than in the meta-analytic approaches. In past work [Hobbs et al. 2012] commensurate priors have been designed to work with patient-level data and covariates.
@@ 2.2 The SPx Prior
A key distinction with RMAP is that SPx uses submodels that are well-separated in the sense that they put most of their prior mass on separate regions of the parameter space for $\left(\theta_{1},\dots,\theta_{H+1}\right)$. This feature is motivated by the literature on nonlocal priors [Johnson and Rossell 2010]. As we later discuss, it leads to posterior inferences about the mechanism and degree of historical borrowing that adapt more quickly to signals in the data. We now describe each component of the SPx prior in detail.
@@
$\theta_{H+1}$ $=m_{hist}\underbrace{\theta_{hist}}_{\text{Expert 1}}+m_{reg}\underbrace{\theta_{reg}}_{\text{Expert 2}}+m_{ind}\underbrace{\theta_{ind}}_{\text{Expert 3}}.$
Here the expert system, or model averaging, prior is obtained by assuming the Categorical prior $(m_{hist},m_{reg},m_{ind})\sim Cat(p_{hist},p_{reg},p_{ind})$. In other words, the $m_{k}$ are binary indicator variables satisfying $m_{hist}+m_{reg}+m_{ind}=1$, and $\theta_{H+1}$ is drawn from model $k\in\{hist,\phantom{,}reg,\phantom{,}ind\}$ with prior probability $p_{k}$. For example, with prior probability $p_{ind}$ the new trial’s logit response rate is $\theta_{H+1}=\theta_{ind}$, which is independent of (completely unrelated to) the historical data. Priors for each expert, i.e. $\theta_{hist},$ $\theta_{reg},\text{ and }\theta_{ind}$, are discussed next.
@@
$w_{h}\propto\left(0.5\right)^{\frac{|\widetilde{\psi}_{h}-\widetilde{\psi}_{H+1}|}{0.05}}$
with $\widetilde{\psi}_{h}:=\text{logit}^{-1}(\bm{\beta}^{T}\mathbf{x}_{h})$. This choice encourages more borrowing from historical trials with covariates closer to $\mathbf{x}_{H+1}$, to the extent that the covariates predict response rates. In particular, the weights are scaled so that historical trial $h$ will receive twice the weight of any trial $j$ for which its predicted response rate is 5% points closer to the predicted response rate of the new trial (i.e. $|\widetilde{\psi}_{j}-\widetilde{\psi}_{H+1}|-|\widetilde{\psi}_{h}-\widetilde{\psi}_{H+1}|=0.05$). We tested other functional forms for the weights and found they typically do not affect posterior inferences for $\hat{\theta}_{H+1}$ greatly. We note that marginally (integrating out just the $\bm{\beta}$ parameter, which has a prior centered on zero), the weights are constant and thus the prior $p\left(\theta_{hist}|\theta_{1},\dots,\theta_{H}\right)$ is centered at the unweighted average $H^{-1}\sum_{h=1}^{H}\theta_{h}$. This is more similar to the commensurate prior, which assumes a constant historical rate across all previous studies (see Section 2.1). While simpler priors may be reasonable in some cases, using a weighted average of potentially varying historical rates with potentially varying weights is a less restrictive modeling assumption.
The prior on $\theta_{reg}$, the regression expert, assumes that the new trial follows the same covariate-response relationship seen in the historical trials. This occurs because the same coefficient $\bm{\beta}$ is used in the regression for $\theta_{h}$ ($h=1,\dots,H$) in (1) and for $\theta_{reg}$. Another key point is that the scale depends on the same $\tau^{2}$ indicating how successful the regression is at predicting the historical response rates, but modified by a constant $c$ so that, after marginalizing out $\tau^{2}$ and $\sigma^{2}$, the prior variance of $\theta_{reg}$ is similar to the prior variance of $\theta_{hist}$ (as we discuss below along with the hyperpriors). At the point $\bm{\beta}=\bm{0}$ the regression expert is equivalent to the classic MAP model. For the regression expert we have assumed a linear regression for the relationship between trial-level covariates and control response rates because the number of historical trials $H$ is typically not large enough to estimate regressions with more complex functional forms. The linear regression may still be useful as a first-order approximation to more complex relationships, and as we discuss in Section 4.4 the use of BMA gives SPx robustness when the linear regression is misspecified.
Lastly, $\theta_{ind}$ is the independent or no-borrowing expert and is included for robustness. We chose the $Logistic(0,1)$ prior for the independent model because it is equivalent to a flat $Uniform(0,1)$ prior on the response probability $\psi_{H+1}$, reflecting the “conservative” analysis that an investigator excluding historical data might choose [Schmidli et al. 2014]. This prior differs only slightly from the $Beta(0.5,0.5)$ prior used by Expert 3 in SPx, which is a Jeffreys prior and may be slightly more efficient [Robert et al. 2009]. As we discuss in Supplementary Section 4, these priors produce nearly identical posterior inferences.
@@
Likelihood: $y_{h}|\mathbf{x}_{h},\psi_{h}$ $\overset{iid}{\sim}Bin(n_{h},\psi_{h}),\hskip 30.00005pth=1,\dots,H,H+1$ (2)
Prior: $\theta_{h}$ :$=\text{logit}(\psi_{h}),\hskip 30.00005pth=1,\dots,H,H+1$
@@
$\theta_{H+1}$ $=m_{hist}\theta_{hist}+m_{reg}\theta_{reg}+m_{ind}\theta_{ind},$
$\theta_{hist}|\mu_{hist},\sigma^{2}$ $\sim N(\mu_{hist},\sigma^{2}),\hskip 30.00005pt\mu_{hist}=\sum_{h=1}^{H}w_{h}\theta_{h},$
$w_{h}$ $\propto\left(0.5\right)^{\frac{|\widetilde{\psi}_{h}-\widetilde{\psi}_{H+1}|}{0.05}},\hskip 30.00005pt\widetilde{\psi}_{h}=\text{logit}^{-1}(\bm{\beta}^{T}\mathbf{x}_{h})$
@@
[…]
$(m_{hist},m_{reg},m_{ind})$ $\sim Cat(p_{hist},p_{reg},p_{ind}).$
@@
The SPx model (2) is completed by hyperpriors for several important parameters. The prior model probabilities $(p_{hist},p_{reg},p_{ind})$ and the hyperpriors for the variances $\sigma^{2}$ and $\tau^{2}$ (and to a lesser degree $\bm{\beta}$) are important because they impact the posterior weighting of the different borrowing mechanisms in SPx. To see how, note that the posterior for $\psi_{H+1}$ is the average of the posteriors under each expert, weighting by the posterior probability that each is “correct”:
$p\left(\psi_{H+1}|D\right)=\sum_{k\in\{hist,reg,ind\}}p(\psi_{k}|m_{k}=1,D)\cdot p(m_{k}=1|D),$ (3)
where $D$ denotes the complete data $(y_{h},n_{h},\mathbf{x}_{h})_{h=1,\dots,H+1}$. The posterior weights of each submodel $k\in\{hist,reg,ind\}$ may be written as
$p^{*}_{k}:=p(m_{k}=1|D)=\frac{p(D|m_{k}=1)\cdot p_{k}}{\sum_{j\in\{hist,reg,ind\}}p(D|m_{j}=1)\cdot p_{j}},$ (4)
where $p(D|m_{k}=1)=\int_{[0,1]^{H+1}}p(D|\psi_{1},\dots,\psi_{H+1})\cdot p(\psi_{1},\dots,\psi_{H+1}|m_{k}=1)\hskip 3.0pt\mathrm{d}^{H+1}(\psi_{1},\dots,\psi_{H+1})$ is the marginal likelihood, or evidence, of expert $k$. It is important to recognize that each expert’s hyperprior affects their marginal prior on the control response rates, $p(\psi_{1},\dots,\psi_{H+1}|m_{k}=1)$. This in turn can greatly influence the expert’s marginal likelihood and posterior model probability, changing the behavior of SPx overall.
@@
where $TCauchy(m,s,I)$ represents a Cauchy distribution with location $m$ and scale $s$ that has been truncated to the interval $I$. Here we use equal prior weights $p_{hist}$ and $p_{reg}$; if an analyst has an a-priori belief that one of these submodels should be trusted more, the prior weights can be made unequal. While $\tau$ has a larger scale than $\sigma$, recall that the (conditional) prior variance of the regression submodel is $c\tau^{2}=\nicefrac{{1}}{{25}}\cdot\tau^{2}$, not $\tau^{2}$, so the regression submodel makes similarly strong prior predictions as the direct historical borrowing one. The prior on the regression coefficients, which does not unduly affect inference, is
@@
following from Gelman et al. [2008].
This choice of prior underlies SPx’s novel strategy to dynamically determine how much historical borrowing is appropriate. Combining (a) and (b) makes SPx more robust to irrelevant historical data as needed in equation (3) primarily by increasing $p^{*}_{ind}$, the posterior probability of the no-borrowing submodel, and only to a much lesser degree by making the posterior inferences $p(\theta_{k}|m_{k}=1,D)$ of the borrowing submodels more conservative. In further detail, (a) makes the borrowing submodels’ marginal likelihoods decrease more quickly as disagreement between the new and historical data grows, and (b) accounts for the fact that the no-borrowing submodel makes less confident predictions overall and thus may have a relatively lower marginal likelihood even when it should be favored. While similar robustness might be sought by giving the borrowing submodels more diffuse priors, this can come at the cost of weaker borrowing when the historical data are actually relevant. We discuss these points further in the supporting information.
@@
Combined with the SPx prior for the new trial’s control rate this induces a prior on the treatment effect, which we define as the difference $\delta:=\psi_{trt}-\psi_{H+1}$ (although other comparisons, such as relative risk, could just as easily be used). Like with other popular methods [Schmidli et al. 2014], in SPx the induced prior for $\delta$ does not have an analytic expression. By construction, the marginal prior for $\delta$ is symmetric with a mean of 0. This is due to the symmetry of the marginal priors for $\psi_{trt}$, $\text{logit}^{-1}\left(\theta_{hist}\right)$, $\text{logit}^{-1}\left(\theta_{reg}\right)$, and $\text{logit}^{-1}\left(\theta_{ind}\right)$, all of which are centered at 0.5 (a 50% response rate). We illustrate the prior for $\delta$ in Figure 1.
Figure 1: Histograms showing the marginal prior distributions for $\psi_{trt}$ (A), $\psi_{H+1}$ (B), and $\delta=\psi_{trt}-\psi_{H+1}$ (C) in the SPx model, based on 10,000 random draws. Priors shown are based on the historical and new trial covariate values used in simulation Scenario 1 (see Section 4).